Hi. Here is my problem:
A cubic box with no top is to be made of plywood. How many square feet of plywood are needed if the box is to hold 64 cubic feet?
Thanks for any help you can give.

Answer by checkley71(8403)

IF (?) THE BOX IS A SQUARE. THEN THE AMOUNT OF PLYWOOD NEEDED FOR A 64 CUBE FOOT BOX IS BASE X~2 WITH A HEIGHT OF X THUS THE VOLUME IS X~2*X =64 OR X~3=64 OR X=4 FEET THUS THEBOX IS 4 FEET BY 4 FEET BY 4 FEET. THEN THE AMOUNT OF PLYWOOD NEEDED IS 4*4 FOR THE BOTTOM & 4*4 FOR EACH OF THE 4 SIDES THUS
4*4+4(4*4)=16+4(16)=16+64=80 SQUARE FEET OF PLYWOOD IS NEEDED.